OpenStudy (anonymous):

p,q,r,s are non zero nos ;p and q are roots of eqn : X^2+rX+s=0;r and s are roots of eqn : X^2+pX+q=0;find p+q+r+s=?

OpenStudy (anonymous):

i have no idea but we can try by example

OpenStudy (anonymous):

damn forgot the minus sign. i should think before i type

OpenStudy (anonymous):

this is what i got \[r+s=-p, p + q = -r\] when you add you get \[r+s+p+q = -p-r\] now i am thinking this means both sides has to be zero, but maybe not. any ideas?